can someone please help me out?
√x3 = 27

3√x2 = 9
i'm totally lost. thank you

Answer by checkley75(3666)

SQRTX^3=27 SQUARE BOTH SIDES WE GET
X^3=27^2
X^3=3*3*3*3*3*3 NOW TAKE THE CUBE ROOT OF X^3 & 9^3
X=9 ANSWER
PROOF
SQRT9^3=27
SQRT(9^2*9)=27
SQRT(9^2*3^2)=27
9*3=27
27=27